OpenStudy (anonymous):

A ball is thrown vertically upward with an initial velocity of 29.4 meters per second. What is the maximum height reached by the ball

OpenStudy (badhi):

using the law of conservation of energy $$\frac{1}{2}mv^2=mgh \implies v=\sqrt{2gh}$$

OpenStudy (anonymous):

\[Hmax = \frac{ u^2 }{ 2g}\]